OpenStudy (anonymous):

Merle's recipe for ketchup calls for sugar, vinegar and tomatoes to be used in a 4: 1: 10 ratio? If she uses 6 cups of sugar, how many cups of vinegar does she needtomatoes should she use?

OpenStudy (anonymous):

for vinegar (6/4) = (x/1) , 6 = 4x = 1.5 cups. for tomatoes (6/4) = (x/10) = 60 = 4x = 15 cups

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i have another problem here.... im really bad at word problems so i appreciate ur help especially with little time left for me..

OpenStudy (anonymous):

sue has a total of 220 points on her first 3 tests. if each test is worth 100 points and there are 10 tests total in the class, how many points will sue have at the end of the course, at this rate. round to nearest point.

OpenStudy (radar):

220:300 :: x:1000 300x=220000 x=733.3